An airplane takes five hours to travel a distance of 3600 km against the wind. The return trip takes four hours with the wind. What is the rate of the plane in still air, and what is the rate of the wind? Thank you for your help!
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

plane speed = x mph
wind speed = y mph
with wind speed = (x+y)
against wind speed = (x-y)
3600/(x+y)=4
4x+4y=3600
/4
x+y=900....................1
3600/(x-y)=5
3600=5x-5y
x-y=720....................2
Add (1) +(2)
2x=1620
x=810 mph speed of plane
900-810= 90 mph speed of wind

Answer by Alan3354(69443)

An airplane takes five hours to travel a distance of 3600 km against the wind. The return trip takes four hours with the wind. What is the rate of the plane in still air, and what is the rate of the wind?
-------------
3600/5 = 720 mph
3600/4 = 900 mph
The airspeed is the average = (720+900)/2 = 810 mph
The windspeed is the difference = 900-810 = 90 mph